OpenStudy (anonymous):

Can Someone Help me Answer This Please Quadratic Equation in One Variable Solve By Factoring 1.4x^2-36=0 2.7x^2+19x=6 3.8x^2+14=-5 4.x^2+8x-9=0 Solve By Quadratic Formula 5.3x^2-2x=8 Solve By Completing Squares 6.x^2-3x-4=0

OpenStudy (dan815):

http://www.youtube.com/watch?v=eF6zYNzlZKQ watch plz! factoring teacher

OpenStudy (anonymous):

1. 4x^2-36=0 collect like terms 4x^2=36,(2x)^2=6^2 find the sqareroot on both sides you get 2x=+or- 6 x=-3,x=3

OpenStudy (anonymous):

do number 3 that same way number 2 and 4,well find the factors

OpenStudy (anonymous):

i slept, sorry my answer in 4 x^2+8x-9=0 (x+9)(x-1)=0 x+9=0 x-1=0 x=9 x=-1

sam (.sam.):

Are your questions graded?

OpenStudy (anonymous):

ye

sam (.sam.):

You are not allowed to post graded questions on openstudy, its against the CoC, because it is considered cheating. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

1.(2x)^2-6^2=(2x+6)*(2x-6) so x=+3,-3 use formula a^2-b^2=(a+b)*(a-b) 2.7x^2+19x-6=0 thn 7x^2+21x-2x-6=0 so x=-3,2/7 4.x^2+9x-x-9=0 thn x(x+9)-1(x+9)=0 so x=-9,1 5.3x^2-2x-8=0 thn use sqt(-2)^2-4*3*8/2*3 so (SQT100)/6 so +10/6 or -10/6